OpenStudy (anonymous):

how many solutions are in this equation? 4x^2+8x+3=0

OpenStudy (anonymous):

(2x+3)(2x+1)=0 x=-3/2 or x=-1/2

OpenStudy (anonymous):

so two solutions? thats not what i asked....

OpenStudy (anonymous):

or 1 ?

OpenStudy (anonymous):

yeah two

OpenStudy (anonymous):

Do you mean..ways how to solve it?

OpenStudy (anonymous):

my bad. its 1. and that is -1/2

OpenStudy (anonymous):

the other one is extraneous